OpenStudy (amberofficial1):

Solve the following problems: 1. x + 7 = 18 2. 9 - y = 14 3. 2x - 4 = 24 4. x - 3 = 7 2

oregonduck (oregonduck):

1) x=11

oregonduck (oregonduck):

3) x=14

oregonduck (oregonduck):

4)x=17

OpenStudy (studytheworld):

1) 11. 2) -5 3)14 4)10

oregonduck (oregonduck):

2) -5

oregonduck (oregonduck):

4 is not 10 dummy

oregonduck (oregonduck):

anyone would know that

OpenStudy (studytheworld):

x-3 = 7. 10-3= 7 No need to be rude.

oregonduck (oregonduck):

x=17 because it would be 17-3=14 14 divided by 2 = 7

OpenStudy (amberofficial1):

THANK YOUU!

OpenStudy (amberofficial1):

Its x-3 over 2

oregonduck (oregonduck):

EXACLTY

oregonduck (oregonduck):

SO... x=17 because it would be 17-3=14 14 divided by 2 = 7

oregonduck (oregonduck):

np need more help?

OpenStudy (mwhit2101):

Calling someone "dummy" is not going to get your point across, or whatever you were attempting to achieve. Please be respectful and polite, not rude and immature @OregonDuck.

OpenStudy (amberofficial1):

please stop fighting. and no im okay thanks for the help.

oregonduck (oregonduck):

i really don't care because in this country we have free speak so if open study is going to punish me for my free speech then there are breaking the are first amendment rights. IN THIS COUNTRY WE HAVE FREE SPEECH RIGHTS SO I CAN SAY ANYTHING I WANT

oregonduck (oregonduck):

so if you punish me or ask someone to do so on openstudy then i can sue you 200,000 for breaking our first amendment rights. I have a lawer and attorney so you try it.

oregonduck (oregonduck):

why do you both medal him in the first place he was wrong and @amberofficial1 you know he is wrong on number 4

OpenStudy (mwhit2101):

Because he was respectful. You're just plain immature. Why are you even on here? You're clearly, like, 14 years old. That's all I'm saying on this subject. Good day.

oregonduck (oregonduck):

what are you going to do @mwhit2101 uhh speech less

oregonduck (oregonduck):

clearly i am older if i know my rights and i am 19

OpenStudy (amberofficial1):

good flutterin bye lmfao

oregonduck (oregonduck):

man how did you not be able you solve problems those were the simplest ones i have soloved in my life you need to learn to do your work be yours if you do it this way you are never going to learn and when you are on to your sat's and act's you are going to do bad. Because you can not solve simple algebra and i go to the usna in annapolis a.k.a the united states of america navy academy

oregonduck (oregonduck):

good day you are a waste of my time